# ReviewAgents: Bridging the Gap Between Human and AI-Generated Paper Reviews

### v1

@@ 2.1 Large Language Models for Paper Review
Large language models (LLMs) have demonstrated significant potential in reviewing and comprehending complex texts Liu and Shah (2023); Zhao et al. (2024); Zhuang et al. (2025); Ruan et al. (2025); Zhu et al. (2024); Ruan et al. (2024b, a). Initial studies, which compared reviews generated by human reviewers and LLMs for academic papers, indicated that LLMs can produce comments with substantial overlap with those of human reviewers, contributing effectively to the peer review process Robertson (2023); Liang et al. (2023). However, further research has shown that, despite the advanced capabilities of LLMs such as GPT-3.5 and GPT-4 in scoring input paper texts, the generated reviews still fall short of fully meeting human expectations Zhou et al. (2024). Researchers have fine-tuned LLMs to accommodate review requirements by constructing datasets from publicly available review comments Kang et al. (2018); Yuan et al. (2021); Shen et al. (2022); Dycke et al. (2023); Gao et al. (2024). Additionally, some studies have employed multi-turn dialogues Tan et al. (2024) or prompted multiple LLMs to construct review systems D’Arcy et al. (2024), providing comprehensive feedback on the entire text. However, unlike current studies that focus solely on generating review comments, our work aims to simulate the entire thought process and review procedure across the full peer review process. We emphasize the individual deep thinking and multi-step, multi-role nature of human peer review in the real world. This approach alleviates biases inherent in reviews generated by a single LLM, thereby aligning more closely with human-generated review feedback.
@@ 2.2 Chain-of-thought Reasoning in Large Language Models
2.3 Instruction Tuning Dataset
Instruction fine-tuning is a specialized training process for LLMs designed to improve their ability to follow specific instructions and perform tasks with higher precision and reliability. This process relies on a dataset containing task-specific examples paired with clear instructions. For academic paper review tasks, various datasets of review comments have been developed, such as those for the NLP community Kang et al. (2018); Dycke et al. (2023) and datasets spanning multiple research domains Gao et al. (2024). Some datasets also include more detailed content, such as meta-review comments Shen et al. (2022) or multi-turn dialogues to demonstrate the rebuttal process Tan et al. (2024). However, these datasets do not capture the reviewer’s thought process. We introduce Review-CoT and ReviewBench, which together establish the first pipeline for LLM-driven automated paper review, encompassing both upstream training and downstream evaluation. To our knowledge, this is the first dataset to structurally represent the review steps.
@@ 3.1 Structured Thinking in Review
Given that paper review requires models to thoroughly read the manuscript, integrate external knowledge, and conduct an in-depth analysis of the paper’s strengths and weaknesses before assigning scores, a detailed stepwise reasoning process is essential. Inspired by the cognitive process of human reviewers, we decompose the review process of the intelligent review agent into three structured stages: Summarization, Analysis, and Conclusion, thereby aligning the automated review procedure with human-like reasoning.
  • Summarization Stage: The agent is provided with the full text of the paper and is tasked with concisely summarizing its core content and methodologies.
  • Analysis Stage: The agent evaluates the strengths and weaknesses of the paper by considering its content in conjunction with relevant works in the same domain. This analysis serves as the foundation for subsequent assessments.
  • Conclusion Stage: Based on the insights gained from the previous two stages, the agent formulates a final review decision for the paper.
@@ 3.3 ReviewBench
Given that the evaluation of LLM-generated review comments is still limited to model scoring and human ratings, with a lack of standardized quantitative assessment criteria, we propose ReviewBench, a benchmark for evaluating LLM-generated review comments across multiple dimensions. Table 2 presents the evaluation dimensions and corresponding metrics included in ReviewBench. Our benchmark encompasses assessments of LLM-generated review comments in terms of language quality, semantic consistency, and sentiment consistency, as well as a Review Arena task to measure the model’s ability to generate review comments. Further details on the evaluation metrics can be found in Appendix.
@@ 5.1 Experimental Setup
We evaluated our method on ReviewBench, selecting advanced open-source models and closed-source models for comparison, evaluating the meta-reviews generated by different models. The evaluation metrics, Distinct and Inverse Self-BLEU, are computed using 4-grams, denoted as $\text{Distinct}_{4}$ and Inverse Self-BLEU@4, respectively. For ROUGE-1 and ROUGE-L, their F1-scores are computed. For the diversity and consistency metrics, including Language Diversity, Semantic Consistency, and Sentiment Consistency, we compute their overall scores.
Our framework utilizes the Llama-3.1-8B-Instruct model trained on NVIDIA A100 GPUs. When generating review comments, we set the number of reviewers, $N$, to 3. The judge for the Review Arena is GPT-4o. Detailed implementation specifics and prompt templates are provided in Appendix.
@@ 5.4 Discussion and Case Study
Figure 4 presents a case study. Compared to comments from advanced LLMs, ReviewAgents’ comments show greater overlap with human reviewers’ comments. For example, both address issues such as insufficient comparison and citation, highlighting the effectiveness of our relevant-paper-aware training approach. While GPT-4o tends to generate more positive comments, ReviewAgents, through stepwise analysis of strengths and weaknesses, produces rejection comments aligned with human reviewers’ assessments. We provide a complete case of the peer review comment generation process in the Appendix.
@@ 7.3 Limited Context Length
The potential risk of our research is the abuse of automatic peer review, which could lead to a decline in the quality of review comments and irresponsible feedback. We emphasize that our work is not intended to replace human reviewers. Instead, they serve as valuable tools for authors to self-edit their papers and as an aid for human reviewers. The results produced by the model should not be misconstrued as definitive and genuine comments on the respective papers.

### v2

@@ 2.1 Large Language Models for Paper Review
Large language models (LLMs) have demonstrated significant potential in reviewing and comprehending complex texts Liu and Shah (2023); Zhao et al. (2024); Zhuang et al. (2025); Zhu et al. (2024). Initial studies, which compared reviews generated by human reviewers and LLMs for academic papers, indicated that LLMs can produce comments with substantial overlap with those of human reviewers, contributing effectively to the peer review process Robertson (2023); Liang et al. (2023). However, further research has shown that, despite the advanced capabilities of LLMs such as GPT-3.5 and GPT-4 in scoring input paper texts, the generated reviews still fall short of fully meeting human expectations Zhou et al. (2024). Researchers have fine-tuned LLMs to accommodate review requirements by constructing datasets from publicly available review comments Kang et al. (2018); Yuan et al. (2021); Shen et al. (2022); Dycke et al. (2023); Gao et al. (2024); Du et al. (2024). Additionally, some studies have employed multi-turn dialogues Tan et al. (2024) or prompted multiple LLMs to construct review systems D’Arcy et al. (2024), providing comprehensive feedback on the entire text. AgentReview Jin et al. (2024) attempts to simulate the interactions among different roles involved in the peer review process, generating review comments and discussion dialogues for each role. However, such interactions do not necessarily occur in real-world review scenarios, and its approach to generating review comments does not adhere to the cognitive patterns typically followed by human reviewers. In contrast to existing work, our approach aims to reconstruct the complete reasoning and evaluation process throughout the peer review workflow. We emphasize the in-depth reflection and multi-step, multi-role nature of real-world peer reviewing, thereby mitigating biases inherent in single-LLM-generated reviews and achieving better alignment with human reviewer expectations.
@@ 3.1 Structured Thinking in Review
Given that paper review requires models to thoroughly read the manuscript, integrate external knowledge, and conduct an in-depth analysis of the paper’s strengths and weaknesses before assigning scores, a detailed stepwise reasoning process is essential. Inspired by the cognitive process of human reviewers, we decompose the review process of the intelligent review agent into three structured stages: Summarization, Analysis, and Conclusion, thereby aligning the automated review procedure with human-like reasoning. In the summarization stage, the agent is provided with the full text of the paper and tasked with concisely summarizing its main contributions and methodologies. During the analysis stage, the agent examines the strengths and weaknesses of the paper based on its content and relevant literature within the same domain, serving as the foundation for subsequent evaluation. In the conclusion stage, drawing upon the outcomes of the previous two stages, the agent formulates a final review verdict for the paper.
@@ 3.3 ReviewBench
Given that the evaluation of LLM-generated review comments is still limited to model scoring and human ratings, with a lack of standardized quantitative assessment criteria, we propose ReviewBench, a benchmark for evaluating LLM-generated review comments across multiple dimensions. Table 2 presents the evaluation dimensions and corresponding metrics included in ReviewBench. Our benchmark encompasses assessments of LLM-generated review comments in terms of language quality, semantic consistency, and sentiment consistency, as well as a Review Arena task to measure the model’s ability to generate review comments. Further details on the evaluation metrics can be found in Appendix A.
@@ 5.1 Experimental Setup
We evaluated our method on ReviewBench, selecting advanced open-source models and closed-source models for comparison, evaluating the meta-reviews generated by different models. The evaluation metrics, Distinct and Inverse Self-BLEU, are computed using 4-grams, denoted as $\text{Distinct}_{4}$ and Inverse Self-BLEU@4, respectively. For ROUGE-1 and ROUGE-L, their F1-scores are computed. For the diversity and consistency metrics, including Language Diversity, Semantic Consistency, and Sentiment Consistency, we compute their overall scores. Detailed implementation specifics and prompt templates are provided in Appendix B and D.
@@ 5.2.2 Review Arena Results
SV CO C CHR PC Average
[…]
Deepseek-V3 3.61 3.85 3.62 3.82 3.92 3.76
GPT-4o 3.80 3.96 4.00 3.84 2.83 3.69
Claude-3.5-sonnet 3.58 3.82 3.78 3.09 3.72 3.60
Qwen-2.5-72B-Chat 3.48 3.79 3.46 2.76 3.11 3.32
GLM-4 2.80 3.57 3.77 3.07 2.75 3.19
[…]
Llama-3.1-8B-Instruct 2.68 3.32 2.85 2.02 2.57 2.69
Table 5: The results of the human evaluation
5.2.3 Human Evaluation Results
We engaged five AI researchers with submission and reviewing experience at top-tier AI conferences such as ICLR, NeurIPS, and *CL to serve as evaluators. They assessed and scored the review comments generated by different LLM-based methods. The evaluation dimensions included Soundness & Validity (SV), Clarity & Organization (CO), Constructiveness (C), Consistency with Human Review (CHR), and Position Clarity (PC). Each dimension was rated on a scale from 1 to 5. Detailed descriptions of each evaluation dimension can be found in Appendix C.
Table 5 presents the results of the human evaluation. ReviewAgents achieved higher average scores compared to proprietary LLMs, with particularly notable improvements in the dimensions of Clarity & Organization (CO) and Consistency with Human Review (CHR). These findings further substantiate that the reviews generated by our approach align more closely with the expectations of human reviewers.
@@ 5.4 Discussion and Case Study
Figure 4 presents a case study. Compared to comments from advanced LLMs, ReviewAgents’ comments show greater overlap with human reviewers’ comments. For example, both address issues such as insufficient comparison and citation, highlighting the effectiveness of our relevant-paper-aware training approach. While GPT-4o tends to generate more positive comments, ReviewAgents, through stepwise analysis of strengths and weaknesses, produces rejection comments aligned with human reviewers’ assessments. We provide a complete case of the peer review comment generation process in the Appendix E.
@@ 7.3 Limited Context Length
The potential risk of our research is the abuse of automatic peer review, which could lead to a decline in the quality of review comments and irresponsible feedback. We emphasize that our work is not intended to replace human reviewers. Instead, our approach aims to enhance the efficiency of the review process while maintaining scientific rigor and alignment with human evaluation standards, rather than replacing human reviewers. The proposed method is designed to serve as an auxiliary tool—integrating LLM-generated reviews as a component of the overall review pipeline—to offer complementary perspectives to traditional expert assessments. In essence, we are exploring how LLMs can augment, rather than supplant, the irreplaceable expertise and discernment of human reviewers. Additionally, our work can serve as a valuable tool for authors to self-edit their papers and as an aid for human reviewers. The results produced by the model should not be misconstrued as definitive and genuine comments on the respective papers.
All annotators were carefully recruited and reasonably compensated for their contributions. They received substantial remuneration to ensure that their time and expertise were fairly valued. Annotators were fully informed of the research objectives and the intended use of the data. All annotators operated under strict institutional guidelines to ensure that data handling adhered to ethical standards, and confidentiality was maintained throughout the annotation and evaluation processes.
Appendix A More Details on ReviewBench and Evaluation Metrics
A.1 Research Areas in ReviewBench
Research Domain # of papers
Machine Learning and Deep Learning 25
NLP and Generative Models 22
Reinforcement Learning 16
Privacy and Security 14
Computer Vision 9
Graph Neural Networks 6
Mathematics and Optimization 5
Application 3
Table 6: The research domains and the number of the papers in ReviewBench
Table 6 presents the research domains of the papers in ReviewBench and their corresponding quantities.
A.2 More Details on Evaluation Metrics
Distinct
Distinct measures the diversity of the generated text by calculating the proportion of unique n-grams in the generated text. A higher Distinct score indicates more variety.
$\text{Distinct}_{n}=\frac{\text{\# of unique n-grams}}{\text{Total \# of n-grams}}$
Self-BLEU
Self-BLEU measures the diversity of the generated text by computing the BLEU score between the generated text and its multiple variants. A higher Self-BLEU score indicates less diversity.
[…]
Where:
  • $r_{i}$ is the i-th variant of the generated text.
  • $\hat{r}_{i}$ is another variant with which the i-th variant is compared.
  • $N$ is the total number of generated texts.
In order to compute the overall score, we employ Inverse Self-BLEU in this paper, where a higher score indicates greater diversity.
ROUGE-1
ROUGE-1 measures the overlap of unigrams (single words) between the generated text and reference text.
Precision=# of matching unigramsTotal # of unigrams in generated $\text{Precision}=\frac{\text{\# of matching unigrams}}{\text{Total \# of unigrams in generated text}}$
Recall=# of matching unigramsTotal # of unigrams in reference $\text{Recall}=\frac{\text{\# of matching unigrams}}{\text{Total \# of unigramsin reference text}}$
[…]
ROUGE-L
ROUGE-L measures the overlap of the longest common subsequence (LCS) between the generated text and the reference text.
Precision=LCS lengthLength of generated $\text{Precision}=\frac{\text{LCS length}}{\text{Length of generated text}}$
Recall=LCS lengthLength of reference $\text{Recall}=\frac{\text{LCS length}}{\text{Length of reference text}}$
[…]
In our work, the calculation of the ROUGE series metrics is done using the rouge_score library in Python.
Appendix B Implementation Details
Our framework utilizes the Llama-3.1-8B-Instruct model trained on NVIDIA A100 GPUs. When generating review comments, we set the number of reviewers, $N$, to 3. The judge for the Review Arena is GPT-4o. We selected advanced open-source models such as Llama-3.1 Dubey et al. (2024), Qwen2.5 Qwen et al. (2025), and closed-source models including GPT-4o, Claude-3.5, and DeepseekLiu et al. (2024); Guo et al. (2025) for comparison. We performed supervised fine-tuning on the base model Llama-3.1-8B-Instruct implemented using Llama-Factory Zheng et al. (2024), setting the number of training epochs to 3, gradient accumulation steps to 32, and warm-up steps to 100. The learning rate was set to 1e-5, with training acceleration implemented using liger-kernel and memory scheduling managed through the Zero3 strategy. The total training cost was approximately 1000 GPU hours.
Appendix C Human Evaluation Questionnaire
Table 7 presents the questionnaire used for human evaluation, detailing the meaning of each evaluation dimension and instructing evaluators to rate the LLM-generated review comments on a scale from 1 to 5. Human-written reviews were also provided alongside as a reference for assessing the relevance and alignment of the generated reviews.
Dimension Description Scoring Guide
Soundness & Validity Does the review provide reasonable, objective evaluations and suggestions based on the content of the paper?
1 = Groundless or incorrect judgments;
5 = Rigorous and well-supported analysis
Clarity & Organization Is the review well-structured, logically organized, and easy to understand?
1 = Disorganized and unclear;
5 = Well-structured and clearly articulated
Constructiveness Does the review offer concrete and valuable suggestions for improvement, rather than just identifying problems?
1 = Vague or unhelpful feedback;
5 = Clear, specific, and helpful feedback
Consistency with Human Review How consistent is the review with real human expert reviews in terms of conclusions, focus points, and suggestions?
1 = Significantly inconsistent;
5 = Highly consistent
Position Clarity Does the review clearly state a stance?
1 = No clear stance or vague position;
5 = Clear and reasonable stance
Table 7: The questionnaire used for human evaluation
Appendix D Prompt Templates
  • Transcription: Figure 10
  • Review generation: Figure 11
  • Meta review generation: Figure 12
  • Review Arena: Figure 13
Appendix E A Case Study of a Complete Process
In Figures 5 to 8, we present the output of a complete review process, providing genuine human review comments for reference (Figure 9). The results indicate that the area chair agent in the ReviewAgents framework effectively synthesizes feedback from different reviewers and generates a well-rounded conclusion. The review conclusions produced by the agent show a high degree of overlap with those made by human reviewers, offering similar feedback on aspects such as the novelty, clarity, lack of theoretical depth, and limited scope of the paper. The final recommendation, similarly, results in an "accept" decision.
Figure 5: Review from reviewer agent 1
Figure 6: Review from reviewer agent 2
Figure 7: Review from reviewer agent 3
Figure 8: Meta review from area chair agent
Figure 9: Meta review from human
Figure 10: Prompt for transcription
Figure 11: Prompt for review generation for Review Agents
Figure 12: Prompt for meta review generation
Figure 13: Prompt used in Review Arena